### v1

# Predictive AI with External Knowledge Infusion for Stocks

Abstract
Fluctuations in stock prices are influenced by a complex interplay of factors that go beyond mere historical data. These factors, themselves influenced by external forces, encompass inter-stock dynamics, broader economic factors, various government policy decisions, outbreaks of wars, etc. Furthermore, all of these factors are dynamic, and exhibit changes over time. In this paper, for the first time, we tackle the forecasting problem under external influence, by proposing learning mechanisms that not only learn from historical trends but also incorporate external knowledge from temporal knowledge graphs. Since there are no such datasets or temporal knowledge graphs available, we study this problem with stock market data, and we construct comprehensive temporal knowledge graph datasets. In our proposed approach, we model relations on external temporal knowledge graphs as events of a Hawkes process on graphs. With extensive experiments, we show that learned dynamic representations effectively rank stocks based on returns across multiple holding periods, outperforming related baselines on relevant metrics.
Machine Learning
1 Introduction
footnotetext: Department of Computer Science and Automation, Indian Institute of Science, Bengaluru, India - 560012.footnotetext: Correspondence to: Ambedkar Dukkipati $<$ambedkar@iisc.ac.in$>$
Figure 1: Depicts the interdependence of stock trends between stocks in the same class. We observe the normalized stock prices of various stocks during the COVID-19 pandemic from January to May 2020 on the NSE. Specifically, after analyzing assets in the pharmaceutical sector, namely CIPLA, DIVISLAB, and DRREDDY, we observed a notable positive trend following the pandemic event. Conversely, assets in the banking industry, such as PNB, INDUSINDBANK, and ICICIBANK, experienced a negative trend during the same period.
One of the most important forecasting problems is related to stock price movements. Contemporary methods model this as time-series forecasting and emply powerful sequence models with deep neural networks.
Existing techniques, such as graph-based embeddings (Li et al., 2019), graph neural networks (Matsunaga et al., 2019), attention mechanisms (Cheng et al., 2018; Qian et al., 2024), event-driven trading (Zhou et al., 2021) and time-aware embeddings (Feng et al., 2019) effectively model the complex relations between the stocks.
Available forecasting solutions in the literature often rely on historical data patterns without considering the external influences that can change with time. However, it is well known that the pattern of evolution of stock prices is not an isolated time-series but a part of a broader financial and economic environment that encompasses various external events, such as government regulations, an outbreak of war, and diseases. These elements introduce a level of unpredictability that makes forecasting challenging, even if one considers the stock interdependencies. Though event-driven trading and decisions based on sentiment analysis of news are well known, there is a notable lack of research on sequential time-series learning that incorporates such external dynamic events.
To study the problem of forecasting in the presence of external influence, we choose the domain of stock markets and propose a new approach that uses temporal knowledge graphs (KG). Considering that no such KGs are available, we construct relevant knowledge graphs and propose representation learning techniques to set benchmark results. Our constructed temporal KGs integrate diverse components, including first and second-order stock relationships, corporate events, macroeconomic indicators, events derived from financial statements, financial ratios disclosed during earnings calls, and analyst sentiments. Notably, all of these elements are equipped with time-aware attributes. The capacity to incorporate entities of varying categories, such as Person, Stock Assets, Products, and Alliance, results in an intrinsically heterogeneous knowledge graph, with these heterogeneous entities acting as mediators or connectors, enabling the modeling of complex second-order relationships in the graph. This constructed graph, named Stock Market Knowledge Graph ($\mathsf{SmKG}$), aims to encapsulate all the pertinent information related to stock assets within a given share market. We use this knowledge graph to construct an algorithm that models market movements along with the occurrence of external events that affect them, resulting in a holistic approach to modeling markets. As part of our method, we model the relations in the temporal knowledge graph as events of a heterogeneous Hawkes process on the graph to learn useful embeddings that aid stock price analysis.
Our key contributions are as follows.
  1. (i)
    We propose a time-series forecasting algorithm that incorporates knowledge of external events influencing the evolution of the series.
  2. (ii)
    For the task of stock prediction, we construct temporal knowledge graphs that capture relevant macroeconomic and financial information related to the NSE and NASDAQ stock markets and use them to enhance sequential forecasting models.
  3. (iii)
    To this end, we propose a time-aware knowledge graph representation technique that leverages a heterogeneous graph convolution operator, along with a temporal point process-based method to model external events. We integrate these embeddings with transformer-based sequential embeddings to obtain accurate rankings of future stock performance over different time horizons.
  4. (iv)
    We empirically demonstrate that our method outperforms related baselines in ranking stocks across three different holding periods. Additionally, we conduct ablation studies and further experiments to analyze the effectiveness of different components of our model.
2 Related Work
Earlier studies (Zhang et al., 2017) on stock movement patterns employ Recurrent Neural Networks (RNNs), relying solely on historical stock price data for analysis, which is inadequate due to ignoring the influence of numerous uncertain political-economic factors prevalent in the real world. In recent times, advancements in automated trading strategies have emerged, exploring event-driven and news-driven trading approaches by detecting relevant events in news articles (Hu et al., 2018; Zhou et al., 2021) and social media texts (Li et al., 2022b). However, these methods assume that stock movements are independent of one another, restricting the ability to model interactions among different stocks. Further, events extracted from raw texts lack the necessary background knowledge pertaining to the entities and relationships they reference.
A new line of research direction focuses on employing graph-based methodologies to represent pairwise relationships between stocks using metadata such as stock-industry information and various connections between companies, such as supply chain associations and links between CEOs and founders. These methodologies tackle stock movement prediction by formulating it as a node classification problem and enhancing traditional historical price-based prediction models through the utilization of graph-based deep learning approaches. Chen et al. (2018) constructs graphs representing shareholder relationships between stocks and utilizes a graph neural network-based model to learn stock movements. Later works (Feng et al., 2019; Matsunaga et al., 2019; Qian et al., 2024) improved the approach by incorporating additional inter-stock relations, such as sector industry connections, supply chain relationships, and inter-executive relationships. They propose a time-aware graph convolutional network that effectively models these inter-stock relations. Ang & Lim (2022) extended this work to incorporate news text data along with inter-stock relations and train their model to generate forecasts of means and volatilities of stock returns over a horizon. Li et al. (2022a) constructs a heterogeneous multi-modal graph that incorporates historical price data, stock news, and technical indicators to model stock trends. Sawhney et al. (2020) and Sawhney et al. (2021) model higher-order inter-stock relationships by formulating them as hypergraphs. However, despite the efficient modeling of different types of data in isolation, there is currently no existing work that comprehensively captures all of these data modalities in a unified framework. Incorporating external knowledge has been studied in the context of natural language processing. Previous works have tackled the issue of grounding natural language to knowledge by aligning or combining language and knowledge graph representations (Wang et al., 2023).
In our method, we model relations on the external dynamic knowledge graph as events of a Hawkes process on a graph, using the learned embeddings for improved predictions. Neural Temporal Point Processes (Mei & Eisner, 2017) have been used in a wide variety of applications to model temporal event streams and predict the occurrence and features of future events. Such processes have further been used to model dynamic interactions between different entities, leading to applications in graph edge prediction and relation forecasting in dynamic graphs (Gracious et al., 2025; Trivedi et al., 2019; Ji et al., 2021). Edge prediction has also been studied without considering them as temporal processes using latent space models (Gracious et al., 2021) and attention mechanisms (Xu et al., 2020).
Figure 2: Different types of relations: (left) without timestamp, (center) with timestamp, and (right) with time interval.
3 Temporal Knowledge Graph Construction
The construction of the knowledge graphs $\mathsf{SmKG}$ involves data acquired from various sources pertaining to the NASDAQ and NSE share markets, mentioned below. Specifically, the stocks listed under the NASDAQ100, S&P500, and NIFTY500 indices are considered. The knowledge graphs are formed to encompass a time span of 20 years, commencing from January 01, 2003, and concluding on December 30, 2022. Statistics of the knowledge graphs for NASDAQ and NSE share markets are shown in Table 1.
Table 1: Statistics of the NASDAQ and NSE $\mathsf{SmKG}$s
$\mathsf{SmKg}$ NASDAQ NSE
# Triples 7,568 2,736
# Quadruples 52,083 44,399
# Quintuples 310,916 251,551
# Entities 4,911 1,049
# Relations 370,567 298,686
# Entity types 12 14
# Relation types 56 53
Figure 3: Visualization of a subgraph from the constructed $\mathsf{SmKG}$s for (left) NASDAQ and (right) NSE share markets. The entities and relations are highlighted based on their types.
General News News articles were acquired by scraping data from two websites, namely “Seeking Alpha” for NASDAQ and “Money Control” for NSE, using the Selenium library in Python [^1] The news articles are obtained by scraping the webpages https://seekingalpha.com/symbol/{stock_ticker}/news for NASDAQ and https://www.moneycontrol.com/company-article/{stock_name}/news/{stock_id} for NSE. A total of 220,645 news articles were collected for NASDAQ and 272,808 for NSE. These articles comprise several components, including the headline, body text, publication time, associated stock asset, related sector, and additional taglines, and contain various types of information such as corporate events, earnings reports, analyst and investor sentiments, etc. To identify the events mentioned in the news articles, a rule-based pattern recognition system was employed for a predefined set of events. The identified events are then added to the constructed knowledge graph $\mathsf{SmKG}$ in the form of quadruple $(h,r,t,[\tau_{s}])$, where $\tau_{s}$ is the news article publication time.
Financial Statements and Ratios The financial statements of companies listed under NASDAQ and NSE were extracted by scraping data from the websites “Macrotrends” and “Top Stock Research” using the Selenium library in Python [^2] The financial statements are obtained from www.macrotrends.net and https://www.topstockresearch.com/ for NASDAQ and NSE respectively.. The collected financial data for each stock consists of various essential components, including the balance sheet, income statement, cash flow statement, and financial ratios. This data is acquired at both annual and quarterly frequencies. Events are derived from the financial statements to identify financial trends by examining whether specific financial parameters exhibit an upward or downward trend. These event trends are then incorporated into the knowledge graph as quintuples, where the interval is defined by the beginning and end of the corresponding financial year or quarter.
Macro-Economic Indicators Macroeconomic indicators provide a snapshot of the overall economic health of a country, encompassing factors such as GDP growth, inflation rate, unemployment rate, trade balance, and fiscal deficit. This information is obtained from “Macrotrends” for both the US and India in which the respective share markets NASDAQ and NSE are located [^3] The macro-economic indicators are obtained from www.macrotrends.net.. The trends in the indicators are identified as events with intervals defined by the beginning and end of the reported year.
Corporate Events Corporate events such as dividend announcements and stock splits are acquired from “Yahoo Finance” for NASDAQ and NSE. These events are represented within the knowledge graph as quadruples, with the timestamp being the date on which the event occurred.
Inter-stock Relationship The inter-stock relationships can be considered to be of two types: first-order and second-order relationships. If there exists a statement with company $i$ as the subject and company $j$ as the object, then they have a first-order relation. Furthermore, companies $i$ and $j$ have a second-order relation if they have statements that share the same object (Feng et al., 2019). For instance, “Apple” and “Foxconn” are in different stages in the production of the product “iPhone”. This data was extracted from “Wikibase,” where the knowledge from “Wikipedia” is stored in the form of triples. A preselected set of relations is identified consisting of $6$ types of first-order relations and $52$ types of second-order relations, which are then filtered from Wikibase.
Sector Classification Each stock asset is classified into a four-level hierarchical system: Sector, Industry group, Industry, and Sub-Industry by Global Industry Classification Standard (GICS) for NASDAQ and Macroeconomic Sector, Sector, Industry, and Basic Industry by NSE Indices for NSE. These relations are added as triples with the asset as head, class as tail, and classification level as relation to the constructed knowledge graph.
4 Proposed Learning Mechanism
Figure 4: TA-HKGE model architecture
4.1 Overview
The input to the model $X^{T}$ at time $T$ consists of $X^{T}=\{\mathcal{P}_{i}^{\tau\in(T-W,T]},G_{\tau_{0}},G_{T}\}$, where $\mathcal{P}_{i}^{\tau\in(T-W,T]}$ is the window of historical stock price data for $W$ previous days of stock $i$ at day $T$, $G_{\tau_{0}}$ is the static subgraph of $\mathsf{SmKG}$, $G_{\tau_{T}}$ is the temporal positive subgraph of $\mathsf{SmKG}$ at time $T$, and $N$ is the total number of stocks considered in the dataset.
The objective is to learn an encoder $\mathcal{E}(X^{T})$ to obtain output embedding $\mathcal{H}^{T}=\left(H_{i}^{T}\right)_{1\leq i\leq N}$. We construct his encoder using multiple components: a Sequential Encoder $\mathcal{E}_{s}(\mathcal{P}_{i}^{\tau\in(T-W,T]})=E_{s}$ of the stock prices, a set of temporal process embeddings $e_{\mathrm{tpp}}$ that are learned to optimally model the events as a heterogeneous Hawkes process, and a Relational Encoder $\mathcal{E}_{r}(G_{\tau_{0}},G_{T},E_{s},e_{\textrm{tpp}})=E_{h}$. These are then passed through the prediction layer $\mathbf{Y}^{\Delta}_{\mathbf{W}_{f}}$ to obtain a vector of ranking scores for the stocks at time step $T+\Delta$.
4.2 Sequential Embedding Layer
Considering the strong temporal dynamics of stock markets, it is intuitive to regard the historical status of a stock as the most influential factor in predicting its future trend. Therefore, we first apply a sequential embedding layer to capture the sequential dependencies in the historical price data and technical indicators. Due to the significant advancements and advantages demonstrated by Transformers in processing sequential data, we choose to utilize a Transformer model to learn the sequential embeddings. Consequently, we feed the historical time series data $\mathcal{P}_{i}^{T}$ of stock $i$ at time step $T$ into the Transformer model, which generates the sequential embedding $s_{i}^{T}$ for each stock as $s_{i}^{T}=\mathsf{Transformer}\left(\mathcal{P}_{i}^{T}\right).$ The Transformer model processes the entire historical sequence simultaneously, capturing long-range dependencies effectively. We employed the encoder-only architecture of the Transformer (Vaswani et al., 2017) to acquire sequential embeddings.
4.3 Temporal Process Embeddings
We then consider relations on the dynamic knowledge graph as events of a heterogeneous Hawkes process on the graph and learn node and relation embeddings using the Heterogeneous Hawkes Process for Dynamic Heterogeneous Graph Embedding (HPGE) algorithm (Ji et al., 2021). For event $(h,r,t,\tau)$, we consider the intensity function $\tilde{\lambda}(h,r,t,\tau)$ defined as
$\mu_{r}(h,t)+\gamma_{1}\sum_{r^{\prime},t^{\prime}\in\mathcal{N}_{r^{\prime}}(h)}m_{1}(h,r,t,\tau,r^{\prime},t^{\prime},\tau^{\prime})$
$\qquad\qquad+\gamma_{2}\sum_{r^{\prime},t^{\prime}\in\mathcal{N}_{r^{\prime}}(t)}m_{2}(h,r,t,\tau,r^{\prime},t^{\prime},\tau^{\prime}),$
where $\mu_{r}$ is the relation type $r$-specific base intensity function between $h$ and $t$, and $m_{1}$ and $m_{2}$ are the mutual intensities between the entities, summing over all relation types $r^{\prime}$ and past $r^{\prime}$-neighborhoods.
Node embeddings $e_{\text{tpp},i}$ and relation embeddings $e^{r}_{\textrm{tpp},j}$ are initialized and the base intensity $\mu_{r}(h,t)$ is taken as
$\left\|\sigma\left(\mathbf{W}_{e}\cdot e_{\text{tpp},h}\right)+e^{r}_{\text{tpp},r}-\sigma\left(\mathbf{W}_{e}\cdot e_{\text{tpp},t}\right)\right\|^{2},$
where $\mathbf{W}_{e}$ is a dense layer and the activation function $\sigma$ is LeakyReLU. The embeddings and weights are optimized by maximizing the intensity function for positive samples and minimizing the intensity function for negative samples. The loss function and the mutual information terms are as defined in Ji et al. (2021).
The obtained node embeddings from the sequential layer and the temporal process model are concatenated to create new node embeddings $E_{s,\text{tpp}}=\left(s_{i}\|e_{\text{tpp},i}\right)_{i}$.
4.4 Relational Embeddings
Figure 5: [Best viewed in color]The experimental setup for model training illustrated across 24 phases. In each phase, the model undergoes a training period of 250 days, followed by a validation period of 50 days and, subsequently, a testing period of 100 days.
To obtain the relational embedding, each entity $i$ in the knowledge graph is assigned a learnable embedding $e^{n}_{i}$ randomly initialized from $\mathcal{N}(0,1)$. The node embeddings $E_{s,\text{tpp}}$ are added to the embeddings of nodes corresponding to the selected stock assets. Together, this creates an entity feature matrix $\mathbf{e}^{n}=\left(e^{n}_{i}\right)_{i=1,\dots,N}$.
Similarly, features for the set of relation types are defined as embeddings $\left(e^{r}_{j}\right)_{j=1,\dots,M}$, where $M$ is the number of relation types, and are initialized from $\mathcal{N}(0,1)$. Each of these is concatenated with the corresponding relation embedding $e^{r}_{\text{tpp},j}$ obtained from the temporal process embedding method and projected to its original dimension to obtain relational embeddings $\mathbf{e}^{r}=\mathbf{W}_{r,\text{tpp}}\cdot\left[e_{r}\|e^{r}_{\text{tpp}}\right]$.
Additionally, embeddings are defined for the month $e^{\tau}_{m}$, day $e^{\tau}_{d}$ and hour $e^{\tau}_{h}$, with weights initialized from $\mathcal{N}(0,1)$. These are then summed together to obtain the temporal embedding for the start timestamp $\tau_{s}$, denoted by $\mathbf{e}^{\tau}$.
These node embeddings $\mathbf{e}^{n}$, along with edge attributes $\mathbf{e}^{e}=\mathbf{e}^{r}+\mathbf{e}^{\tau}$ are passed through two layers of Heterogeneous Edge-featured Graph Attention (HEAT) (Mo et al., 2021) and two linear layers to obtain the final combined relational embedding $E_{h}$ for all stocks.
4.5 Prediction Layer
The obtained embeddings are passed through a prediction layer to obtain a vector of ranking score
$\hat{Y}^{T+\Delta}=\frac{\exp(\mathbf{W}_{f}(E_{s,\text{tpp}}\|E_{h})+b_{f})}{\sum_{n=1}^{N}\exp(\mathbf{W}_{f}(E_{s,\text{tpp}}\|E_{h})+b_{f})}$ (1)
that acts as the predicted rank of the given stocks in the non-increasing order of their potential return on investment from time $T$ to $T+\Delta$.
4.6 Loss Function
The loss function is defined as the sum of four terms as follows.
$\mathcal{L}$ $=\alpha_{1}\mathcal{L}_{1}(h,r,t,\tau)+\alpha_{2}\mathcal{L}_{2}(\{\hat{Y}\},\{Y\})$
$\qquad+\alpha_{3}\mathcal{L}_{3}(\{\hat{Y}\},\{D\})+\alpha_{4}\mathcal{L}_{4}(\{\hat{Y}\},\{R\}).$ (2)
Here $\mathcal{L}_{1}(h,r,t,\tau)$ is the Knowledge Graph Embedding loss, which ensures consistency of the learned node and temporal embeddings by minimizing
$\mathcal{L}_{1}(h,r,t,\tau)=\|e^{n}_{h}+e^{r}+e^{\tau}-e^{n}_{t}\|,$
$\mathcal{L}_{2}$ is the listwise loss function ApproxNDCG (Bruch et al., 2019) comparing the actual and predicted ranks, $\mathcal{L}_{3}$ is a binary cross-entropy loss between the prediction and an indicator variable whether the stock price moved in the upward direction, and finally, $\mathcal{L}_{4}$ is another cross-entropy loss between the predicted ranking score and an indicator whether the stock is in the top 5 with respect to the IRR metric. The detailed expressions for each of the above losses are provided in Appendix A.
5 Experimental Setup
Table 2: Statistics of the three datasets - NASDAQ100, S&P500 and NIFTY500.
- NASDAQ100 S&P500 NIFTY500
#Assets 83 442 336
#Days 2800 2800 2800
Table 3: Comparison of the obtained metrics of TA-HKGE with the baseline models. The datasets NASDAQ100, S&P500, and NIFTY500 are represented as N100, S&P, and NIFTY. The best result for each dataset and holding period is highlighted in bold. It can be noted that TA-HKGE performed better in most metrics.
$\Delta$ Model Dataset Top 1 Top 5
IRR AIRR SR NDCG ACC IRR AIRR SR NDCG ACC
1 TRANSF N100 0.088 25.027 0.874 0.214 1.014 0.08 22.498 0.09 0.315 5.556
S&P 0.062 16.868 0.748 0.143 0.181 0.065 17.9 0.03 0.208 1.103
NIFTY 0.108 33.477 0.484 0.152 0.389 0.098 28.173 -1.02 0.222 1.372
STHGCN N100 0.054 14.654 0.761 0.209 0.861 0.082 22.971 0.117 0.309 4.931
S&P 0.054 14.841 0.667 0.139 0.194 0.052 13.969 -0.02 0.204 0.8
NIFTY 0.055 14.982 0.62 0.21 0.82 0.069 19.114 0.06 0.31 4.74
STHAN N100 0.055 14.982 0.617 0.209 0.819 0.069 19.114 0.057 0.308 4.744
S&P 0.056 15.258 0.589 0.14 0.153 0.059 16.013 0.007 0.205 0.914
NIFTY 0.087 24.658 -1.32 0.15 0.347 0.112 32.582 -0.72 0.22 1.433
GCNKG N100 0.09 27.302 0.88 0.213 1.056 0.1 28.741 0.135 0.312 5.136
S&P 0.10 27.84 0.92 0.142 0.292 0.07 19.24 0.02 0.207 1.072
NIFTY 0.128 38.14 -0.621 0.148 0.278 0.109 31.538 -1.08 0.218 1.233
TA-HKGE N100 0.09 25.44 0.18 0.291 5.00 0.11 31.92 0.1 0.397 13.15
S&P 0.05 13.42 0.376 0.163 1.58 0.08 22.32 0.059 0.228 2.30
NIFTY 0.2 64.44 2.32 0.18 0.25 0.17 53.42 0.05 0.246 4.63
$\Delta$ Model Dataset Top 1 Top 5
IRR AIRR SR NDCG ACC IRR AIRR SR NDCG ACC
5 TRANSF N100 0.403 22.524 1.787 0.214 1.125 0.422 23.675 0.551 0.316 5.736
S&P 0.247 13.263 1.307 0.139 0.042 0.276 14.882 0.42 0.206 0.942
NIFTY 0.457 25.848 1.36 0.147 0.25 0.501 28.633 0.071 0.22 1.317
STHGCN N100 0.455 25.744 1.926 0.215 1.306 0.484 27.559 0.623 0.318 6.114
S&P 0.323 17.64 1.542 0.142 0.306 0.328 17.961 0.482 0.208 1.242
NIFTY 0.498 28.956 1.149 0.149 0.472 0.502 28.737 0.049 0.22 1.411
STHAN N100 0.466 26.507 2.09 0.214 1.236 0.489 27.917 0.611 0.318 6.197
S&P 0.45 25.6 2.08 0.144 0.5 0.363 20.028 0.517 0.209 1.217
NIFTY 0.564 33.529 -1.34 0.145 0.292 0.489 27.867 0.031 0.217 1.322
GCNKG N100 0.414 23.165 2.015 0.211 0.833 0.417 23.333 0.545 0.316 5.744
S&P 0.395 22.119 2.029 0.14 0.153 0.308 16.801 0.48 0.206 1.025
NIFTY 0.59 35.72 1.44 0.148 0.347 0.516 29.757 0.065 0.219 1.356
TA-HKGE N100 0.74 45.0 0.853 0.296 6.58 0.62 36.54 0.597 0.397 13.33
S&P 0.37 20.45 1.15 0.166 1.5 0.18 9.48 0.20 0.222 2.9
NIFTY 0.8 49.42 2.29 0.153 0.37 0.65 38.61 0.652 0.227 2.13
$\Delta$ Model Dataset Top 1 Top 5
IRR AIRR SR NDCG ACC IRR AIRR SR NDCG ACC
20 TRANSF N100 1.641 22.782 3.78 0.21 1.264 1.711 23.829 1.382 0.315 5.95
S&P 1.029 13.769 2.892 0.139 0.153 1.126 15.154 1.142 0.205 1.239
NIFTY 2.152 30.789 3.369 0.147 0.264 2.31 33.351 1.225 0.221 1.442
STHGCN N100 1.702 23.719 4.37 0.207 1.069 1.718 23.935 1.371 0.314 6.2
S&P 1.685 23.617 3.964 0.143 0.653 1.49 20.46 1.34 0.209 1.619
NIFTY 2.633 38.794 4.092 0.15 0.92 2.358 34.141 1.28 0.221 1.586
STHAN N100 1.742 24.328 4.338 0.208 0.931 2.013 28.548 1.54 0.316 6.819
S&P 1.79 25.13 4.2 0.144 0.847 1.376 18.846 1.165 0.211 1.567
NIFTY 2.95 44.49 4.52 0.147 0.486 2.388 34.657 1.33 0.22 1.592
GCNKG N100 1.504 21.345 3.848 0.211 1.417 1.404 19.654 1.126 0.319 6.372
S&P 1.172 16.001 3.196 0.138 0.181 1.168 15.794 1.085 0.206 1.253
NIFTY 2.632 38.978 4.109 0.147 0.569 2.47 36.06 1.294 0.22 1.575
TA-HKGE N100 3.51 54.46 6.067 0.289 4.87 2.0 28.34 0.718 0.391 12.99
S&P 1.5 20.63 4.4 0.166 1.08 1.56 21.53 0.907 0.235 3.53
NIFTY 6.65 125.06 6.585 0.150 1.42 1.55 21.38 1.88 0.236 2.67
Three datasets are created containing a subset of stocks from NASDAQ and NIFTY, with datasets NASDAQ100 and S&P500 from the NASDAQ market and NIFTY500 from NSE, using Yahoo Finance. Each dataset consists of historical price movements and the constructed knowledge graph valid for each trading day. The historical price movements include features Close, Open, Low and High prices, and trading volume for each trading day from 30 January 2012 to 20 October 2022. The statistics are given in Table 2.
The model is given as input a context window of stock data $X_{T}$ at time $T$, which consists of the historical stock price $\mathcal{P}^{\tau\in(T-W,T]}$ and the positive subgraph $G_{T}$ from the knowledge graph at the time point $T$. To avoid data leakage and minimize distribution drift, the values in the historical price window are normalized by dividing them by the last value in the previous window.
For learning embeddings using the heterogeneous Hawkes process, we transformed the data tuples into the format (head, head_type, tail, tail_type, relation, timestamp), with a separate tuple generated for each month within the range of validity of the relation.
Table 4: TPP Shows the performance of the TA-HKGE model with a 750-day training period (750TR) and a 1000-day training period (1000TR).
$\Delta$ Model Dataset Top 1 Top 5
IRR AIRR SR NDCG ACC IRR AIRR SR NDCG ACC
1 750TR N100 0.112 32.47 0.652 0.268 2.5 0.121 35.67 0.428 0.375 12.12
S&P 0.004 0.92 0.179 0.154 0.25 0.066 18.02 0.306 0.225 2.62
NIFTY 0.139 41.83 -499.2 0.163 0.57 0.094 26.7 0.470 0.233 1.97
1000TR N100 0.059 16.18 0.266 0.271 0.377 0.123 36.22 0.448 0.382 11.58
S&P 0.158 48.78 0.463 0.162 1.16 0.078 21.81 0.329 0.233 2.15
NIFTY 0.087 24.38 0.515 0.155 0.17 0.111 32.4 0.485 0.227 1.83
5 750TR N100 0.38 21.58 1.043 0.275 2.49 0.617 36.34 1.035 0.384 12.46
S&P 0.718 43.42 1.923 0.166 1.05 0.457 25.85 0.856 0.236 3.39
NIFTY 0.447 25.2 1.008 0.159 0.5 0.332 18.21 0.665 0.231 1.93
1000TR N100 0.411 22.97 0.984 0.279 3.27 0.645 38.25 1.006 0.390 13.81
S&P 0.142 7.4 0.597 0.169 1.03 0.621 36.65 1.433 0.244 4.14
NIFTY 0.901 57.1 2.102 0.157 0.71 0.638 37.76 1.181 0.23 2.16
20 750TR N100 0.378 4.87 1.536 0.258 2.17 1.456 19.98 1.698 0.3723 11.58
S&P 0.588 7.66 1.254 0.187 1.35 1.256 17.03 1.420 0.258 4.34
NIFTY 2.637 38.8 3.404 0.162 0.1 2.332 33.7 2.107 0.235 2.51
1000TR N100 1.334 18.18 1.433 0.265 4.01 2.082 29.64 2.440 0.373 11.35
S&P 0.907 12.05 1.767 0.162 0.11 1.853 26.02 1.945 0.237 3.6
NIFTY 0.613 8.0 0.565 0.147 0 1.679 23.34 1.724 0.219 1.33
5.1 Training Setup
The value of $\Delta$, the number of days in the future to be considered for prediction, is set as 1, 5, or 20, corresponding to daily, weekly, or monthly holding periods, respectively. As illustrated in Figure 5, we divide the dataset into 24 smaller overlapping datasets to train and evaluate the model separately in different phases by shifting a 600-day sliding window over the entire trading period, starting from a 400-day window. Each phase is divided into a training period of 450 trading days, a validation period of 50 days, and a testing period of 100 days, with the first few phases starting from a 250-day training period. The calculated metrics are obtained by averaging the results from all the test sets across the 24 phases. The reported metrics are described in Appendix B.
5.2 Hyperparameters
Table 5: Hyperparameters
Parameter Value
Learning Rate 0.0006
Batch size 128
Number of epochs 10
Temporal Point Process embeddings size 128
Learning rate for HPGE model 0.0001
Important hyperparameters used for the algorithm are mentioned in the Table 5. The weight coefficients $\alpha_{i}$ for each loss component in the objective are set to $1$. We performed experiments varying the size of the Relational Embedding in $\{16,32,64,128\}$ and chose the best performing dimension. We set the size of the sequential embeddings to 20 and that of the relation embeddings to 16 for NIFTY and 128 for the NASDAQ and S&P. For ablation studies, we used embedding sizes of 20 and 16 respectively to evaluate their effect on model performance.
Table 6: TPP Standard deviation of metrics across all 24 phases using our model TA-HKGE. The standard deviation is lower when more stocks are selected.
$\Delta$ Dataset Top 1 Top 5
NDCG ACC IRR SR NDCG ACC IRR SR
1 N100 0.037 3.24 0.27 1.62 0.032 3.22 0.19 0.35
S&P 0.029 2.58 0.23 1.76 0.023 1.6 0.15 0.41
NIFTY 0.008 0.49 0.30 2.86 0.006 0.84 0.17 0.41
5 N100 0.05 6.40 1.42 4.23 0.032 3.94 0.84 0.69
S&P 0.022 2.39 1.11 3.37 0.018 1.42 0.85 1.07
NIFTY 0.023 1.79 1.36 3.42 0.02 2.24 0.96 0.86
20 N100 0.05 4.66 5.63 8.26 0.047 5.39 4.25 3.08
S&P 0.035 3.01 3.42 13.89 0.032 3.27 2.25 1.94
NIFTY 0.019 1.41 6.34 11.56 0.05 2.16 3.98 2.53
Table 7: Comparison of TA-HKGE without various components of the architecture as described in Section 5.4.
$\Delta$ Model Dataset Top 1 Top 5
IRR AIRR SR NDCG ACC IRR AIRR SR NDCG ACC
1 WOSEQ N100 0.13 38.73 0.60 0.27 3.87 0.08 22.32 0.04 0.37 10.82
S&P 0.103 19.61 0.77 0.15 0.49 0.08 22.32 0.05 0.221 1.95
NIFTY 0.15 45.89 0.85 0.154 0.29 0.14 42.26 0.244 0.225 1.67
WOHK N100 0.13 38.73 0.61 0.29 5.24 0.13 38.73 0.14 0.40 14.4
S&P 0.59 340.36 0.23 0.17 1.37 0.10 28.64 0.078 0.24 3.44
NIFTY 0.06 16.31 0.425 0.157 0.33 0.13 38.73 0.22 0.22 2.08
LSTM* N100 0.16 49.61 0.77 0.28 5.12 0.13 38.73 0.14 0.39 12.54
S&P 0.016 4.11 0.279 0.16 0.62 0.06 16.31 0.031 0.23 2.57
NIFTY 0.17 53.42 1.00 0.15 0.58 0.12 35.28 0.17 0.22 1.76
WOTPP N100 0.09 25.44 0.48 0.27 3.58 0.11 31.92 0.11 0.37 10.77
S&P 0.11 31.92 0.90 0.16 0.79 0.09 25.44 0.99 0.22 2.23
NIFTY 0.18 57.33 1.14 0.156 0.29 0.13 38.73 0.19 0.23 1.65
5 WOSEQ N100 0.55 31.84 1.38 0.27 3.87 0.58 33.84 0.54 0.38 13.39
S&P 0.06 3.06 0.56 0.16 1.20 0.236 12.61 0.34 0.22 2.36
NIFTY 0.55 31.84 0.86 0.151 0.24 0.39 21.67 0.53 0.227 1.56
WOHK N100 0.50 28.57 0.23 0.31 6.02 0.68 40.71 0.57 0.41 15.8
S&P 0.624 36.82 1.90 0.17 1.66 0.417 23.33 0.49 0.24 4.14
NIFTY 0.745 45.36 1.86 0.16 0.79 0.63 37.23 0.646 0.23 2.65
LSTM* N100 0.68 40.71 1.51 0.28 4.83 0.5 28.57 0.49 0.38 12.6
S&P 0.68 40.71 1.64 0.17 2.08 0.36 19.85 0.36 0.24 3.74
NIFTY 0.72 43.55 1.45 0.15 0.99 0.72 43.55 0.76 0.22 2.10
WOTPP N100 0.71 42.84 2.18 0.28 4.24 0.6 35.18 0.586 0.38 12.7
S&P 0.77 47.19 3.07 0.16 0.95 0.55 31.84 0.68 0.22 2.51
NIFTY 0.98 63.47 2.32 0.15 0.54 0.62 36.54 0.56 0.22 1.61
20 WOSEQ N100 3.35 51.46 5.22 0.29 7.83 1.92 27.07 0.86 0.39 13.36
S&P 1.68 23.35 0.45 0.177 2.37 1.36 18.55 0.89 0.24 4.10
NIFTY 3.53 54.82 2.57 0.153 0.66 2.48 36.16 1.37 0.229 2.07
WOHK N100 0.84 36.16 0.617 0.28 3.33 1.43 19.59 0.616 0.39 13.98
S&P 0.14 1.77 -3.11 0.179 1.87 0.92 12.23 0.16 0.24 4.77
NIFTY 2.48 36.16 2.96 0.158 0.24 2.89 43.186 1.84 0.23 3.52
LSTM* N100 3.66 57.29 3.13 0.30 5.74 2.18 31.22 0.38 0.40 15.13
S&P 0.38 4.89 -0.84 0.18 1.16 1.05 14.06 0.255 0.24 4.59
NIFTY 2.08 29.61 1.81 0.15 0.62 2.64 38.86 1.24 0.23 2.18
WOTPP N100 2.43 35.32 3.48 0.28 4.58 1.42 19.44 0.68 0.379 11.34
S&P 1.95 27.54 3.20 0.182 2.20 1.40 19.14 0.738 0.24 4.24
NIFTY 6.09 110.61 6.96 0.16 1.66 2.22 31.87 1.54 0.24 2.89
5.3 Baselines
First, we consider a simple (TRANSF) (Vaswani et al., 2017) acting on the historical price information $\mathcal{P}$. Next, we consider the Spatio-Temporal Hypergraph Convolution Network (STHGCN) (Sawhney et al., 2020), which uses a gated temporal convolution over hypergraphs constructed from the stock asset sector classifications, and the Spatio-Temporal Hypergraph Attention Network (STHAN-SR) (Sawhney et al., 2021), which models stock dependencies via hypergraphs to capture higher-order relations, with a hyperedge indicating the presence of a first-order or second-order relation. Finally, we consider GCNKG, a two-layer GCN that considers a homogenous directed graph version of the current time snapshot of the constructed knowledge graphs to obtain node embeddings. These are then concatenated with sequential embeddings and used to calculate the prediction scores. The loss function used for training the baselines is the linear combination of pair-wise ranking loss (Equation 4) and direction loss (Equation 7) defined in Appendix B. All the hyperparameters used for training the baseline and our model are the same.
5.4 Ablations and Additional Experiments
How much does the sequential embedding as asset features and modeling heterogeneous entities help in model performance?
We wish to study how much the components of our method affect the model performance. To that end, we empirically compare multiple variants of our model TA-HKGE without Temporal point process. For the first ablation WOTPP, we remove the embeddings obtained from the temporal process modeling and only consider sequential and relational embeddings. For the second variant, LSTM*, we replace the transformer-based sequential embedding in WOTPP with an LSTM. Next, we consider WOSEQ, in which the sequential embedding update for the stock assets as input for the relational embedding layer is removed, along with the embeddings obtained from the temporal point process model. Finally, the temporal point process embeddings and the heterogeneous transformation of entity features are removed to obtain WOHK. Table 7 provides detailed results, from which it is clear that each of the components of our architecture plays an important role in improving the performance of the algorithm.
How well the model perform when the training, validation, and testing time in each phase is increased?
It is also essential to see how well the model with temporal point process performs when the training, validation, and testing times of each phase are increased since drifts in the stock price over time and changes in the nature of external events can have a negative effect on the model. To measure the extent of this effect, two more training setups are considered. In the first setup ($750$TR), we consider a training period of 750 trading days (3 years) with 75-day validation and a 200-day testing period. In the next setup (1000TR), the training period is 1000 trading days (4 years), with 100-day validation and a 300-day testing period. Table 4 shows that the performance is affected when the training period is increased. However, the results obtained from the model still beat risk-free returns as indicated by the Sharpe ratio.
Table 8: Best and Worst Case ROI Scenarios
$\Delta$ Dataset Top 1 (IRR) Top 5 (IRR)
Best Worst Best Worst
1 N100 4.06 -3.65 3.67 -3.26
S&P 2.98 -2.79 2.65 -2.41
NIFTY 5.12 -3.41 3.77 -3.43
5 N100 4.60 -3.10 4.14 -2.89
S&P 3.62 -2.66 2.78 -2.27
NIFTY 4.69 -2.91 4.52 -3.09
20 N100 4.26 -3.22 4.80 -2.21
S&P 4.68 -1.58 4.33 -1.12
NIFTY 6.13 -2.56 5.61 -2.32
What is the best and worst case scenario in terms of return on investment?
Since, in practice, it is a difficult task to find the exact value of the closing price of a stock, the best and worst-case scenario in terms of IRR is calculated as well. The best case scenario is when the stock is bought at its low price on day $T$ and sold at its high price on day $T+\Delta$, whereas the worst case is when the stock is bought at its high price on day $T$ and sold at its low price on day $T+\Delta$. Table 8 shows the obtained best and worst case IRR using our model, including the Temporal point process. It is observed that the variance in the best and worst IRR decreases as the holding period increases. Also, it is to be noted that accurately predicting the closing price is also an important task for real-life applications of this work.
What is the standard deviation of metrics across the 24 phases?
The market can be bullish or bearish during certain phases, so it is important to look at the standard deviation of the obtained metrics across all the phases. Table 6 shows the standard deviation of metrics across all 24 phases obtained using our model with a Temporal point process. The standard deviation is comparatively lower for the larger datasets S&P500 and NIFTY500. Also, it can be noticed that the standard deviation is lower when the number of stocks selected is higher.
6 Conclusion
In this paper, for the first time, we try to address the problem of forecasting under external influence. To set benchmarks we built Stock Market Knowledge Graphs ($\mathsf{SmKG}$) that offer a systematic framework for constructing knowledge graphs encompassing comprehensive information about publicly available share market data. Our knowledge graphs are designed to capture the temporal and heterogeneous nature of the data and are specifically constructed for two real-world share markets, namely NASDAQ and NIFTY. By integrating various sources of information, including general news events, financial statements and ratios, macroeconomic indicators, corporate events, inter-stock relationships, and sector classification, our framework provides a unified and consolidated view of the share market.
Furthermore, we introduce a model called TA-HKGE, which leverages the power of heterogeneous graph attention networks, temporal process models, and time-aware translation-based knowledge graph embedding techniques to effectively learn the underlying structures and relationships driving stock prices using the constructed knowledge graph. We empirically demonstrate the validity and effectiveness of our proposed methods on the task of stock ranking based on relevant metrics and showcase the importance and necessity of the various components incorporated in our framework. Overall, our research contributes to the field of share market analysis by providing a comprehensive and efficient framework to solve a forecasting problem by incorporating external knowledge in the form of a temporal knowledge graph.
Appendix
Appendix A Loss Functions
A.1 KGE Loss Function
The time-aware knowledge graph embeddings are learned using a translation-based model type loss function, in which the element-wise addition of head entity, relation embedding, and temporal embedding should be close to the tail embedding.
$\mathcal{L}_{1}(h,r,t,\tau)=\|e^{n}_{h}+e^{r}+e^{\tau}-e^{n}_{t}\|$ (3)
A.2 Listwise Ranking
Earlier works (Feng et al., 2019; Sawhney et al., 2020, 2021) employ the following pairwise loss function as the objective.
$\sum_{i=0}^{N}\sum_{j=0}^{N}\max\left\{0,-\left(\hat{Y}_{i}^{\tau+\Delta}-\hat{Y}_{j}^{\tau+\Delta}\right)\left(Y_{i}^{\tau+\Delta}-Y_{j}^{\tau+\Delta}\right)\right\},$ (4)
where $\hat{Y}_{i}^{\tau+\Delta}$ is the predicted ranking score and $Y_{i}^{\tau+\Delta}$ is the actual ranking score for stock $i$. This approach computes the loss for each pair of scores separately, which might be too loose as an approximation of the performance measures of NDCG and MAP. By contrast, the listwise loss function ApproxNDCG (Bruch et al., 2019) used in the listwise approach can better represent the performance measures. The listwise ranking loss function is given by,
$\mathcal{L}_{2}(\{\hat{Y}\},\{Y\})$ $=-\frac{1}{\text{DCG}(\hat{y},\hat{y})}\sum_{i}\frac{2^{\hat{y}_{i}}-1}{\log_{2}(1+\text{rank}_{i})},$ (5)
$\text{where ~{}rank}_{i}$ $=1+\sum_{j\neq i}\frac{1}{1+\exp\left(\frac{-(y_{j}-y_{i})}{\text{temperature}}\right)}.$ (6)
Here, $y$ is the predicted ranking score and $s$ is the actual score, which is the IRR for each stock from day $\tau$ to day $\tau+\Delta$.
A.3 Cross-Entropy Losses
Let $D^{\tau+\Delta}_{i}$ be the binary movement direction from the closing price of stock $i$ from day $\tau$ to $\tau+\Delta$, in which 1 denotes rise, and 0 denotes fall, and it is expressed as $D^{\tau+\Delta}_{i}=1(p^{i}_{\tau+\Delta}\geq p^{i}_{\tau})$ where, $p^{i}_{\tau}$ is the closing price of stock $i$ on the day $\tau$. The movement ranking loss is defined as the binary cross entropy between $\hat{Y}^{\tau+\Delta}_{i}$ and $D^{\tau+\Delta}_{i}$.
$\mathcal{L}_{3}(\{\hat{Y}\},\{D\})=-[D\log(\hat{Y})+(1-D)\log(1-\hat{Y})]$ (7)
Also, another cross-entropy loss is calculated between the predicted ranking score and stocks that appeared in the top $K$ rank. Such that $R^{\tau+\Delta}_{i}=1(IRR_{i}\text{is in Top 5})$ is the truth labels.
$\mathcal{L}_{4}(\{\hat{Y}\},\{R\})=-[R\log(\hat{Y})+(1-R)\log(1-\hat{Y})]$ (8)
Appendix B Evaluation Metrics
B.1 Investment Return Ratio (IRR)
The Cumulative Investment Return Ratio (IRR) is defined as:
IRR=Final ValueInitial ValueInitial $\text{IRR}=\frac{\text{Final Value}-\text{Initial Value}}{\text{Initial Value}}\times 100,$
the percentage difference between the value of the investment after and at the start of the holding period.
B.2 Annualized Investment Return Ratio (AIRR)
The Annualized Investment Return Ratio (AIRR) is defined as
$\text{AIRR}=(1+\text{ROI})^{\frac{252}{\Delta}}-1,$ (9)
where $252$ is the average number of trading days in a year, $\Delta$ is the number of holding days, and ROI is the Return On Investment defined as IRR$/100$. The AIRR quantifies the average annual rate of return achieved on the investment, taking into account the compounding effect over time.
B.3 Sharpe Ratio (SR)
The Sharpe ratio measures the risk-adjusted return of an investment and can be calculated using the following formula:
SR $=\frac{R_{p}-R_{f}}{\sigma_{p}}$ (10)
where, $R_{p}$ represents the average return of the investment, $R_{f}$ represents the risk-free rate of return, and $\sigma_{p}$ represents the standard deviation of the investment’s returns. The Sharpe ratio quantifies the excess return earned per unit of risk taken. It is commonly used to assess the attractiveness of an investment opportunity based on its risk-adjusted performance. The risk-free rates are calculated for each phase by averaging the 6-month treasury-bill yield obtained from “www.treasurydirect.gov” for the USA and “www.investing.com” for India.
B.4 Normalized Discounted Cumulative Gain (NDCG)
The Discounted Cumulative Gain (DCG) is a metric used to evaluate the quality of a ranked list. It can be calculated as
[…]
where, $\text{rel}_{i}$ represents the relevance score of the item at position $i$ in the ranked list. The DCG assigns higher weights to more relevant items at the top of the list by dividing the relevance score by the logarithm of the position. The DCG accumulates these weighted scores up to position $k$ to measure the overall quality of the ranked list. The Ideal Discounted Cumulative Gain (IDCG) is a reference value used to evaluate the quality of a ranked list, where $\text{rel}_{i}$ represents the relevance score of the item at position $i$ in the ideal ranking order. The Normalized Discounted Cumulative Gain (NDCG) is a metric used to evaluate the quality of a ranked list. It can be calculated using the following formula:
[…]
The NDCG provides a normalized measure of the ranking performance, ranging from 0 to 1, with higher values indicating better ranking quality.
B.5 Accuracy (ACC)
The accuracy represents the ratio of correct predictions to the total number of predictions made by the model. It provides an overall measure of the model’s ability to correctly identify the top $k$ stocks at each time point.
[…]